Use the order of operations to simplify each expression
A 36-24/4.3-1
b -8(-4 -5 (-3)

Please help solve this

Answer by tanimachatterjee(60)

36-24/4.3-1
= 36 - 6.3 -1
= 36 - 18 -1
= 36 -( 18 +1)
= 36 - 19
= 17
b -8(-4 -5 (-3)[one closing bracket is missing]
= -8(-4 +15)
= -8(11)
= -88